-x+2y=7
Y=-2x+1
How would I solve this?

Answer by hamsanash1981@gmail.com(151)

-x + 2y = 7 --- 1
-2x -y = -1 ---2
Multiplying equ 2 by 2
-4x - 2y = -2 --- 3
Solving equation 1 and 3
-x + 2y = 7
-4x -2y = -2
=============
-5x = 5
x = -1

Substituting value of x in equ 1
1 + 2y = 7
2y = 6
Y=3
Values of X= -1 and Y= 3